# FEUDA: Frustratingly Easy Prompt Based Unsupervised Domain Adaptation

### v1

@@ 7 Conclusion
We introduce the setting of prompt-based UDA, where the discriminative prediction is converted into a generative task. We then study the necessity of continued pre-training and domain-invariance based methods for UDA by introducing FEUDA, a “frustratingly easy” prompt-based UDA method. FEUDA involves training an auto-regressive LM on the unlabeled source and target data through the MLM task, followed by supervised training on the labeled source data. Across various datasets, models, adaptation methods, and few-shot settings, FEUDA is competitive with strong UDA methods that promote domain invariance. We also investigate the impact of continued pre-training on the UDA setup. We discover that the MLM task aids the model in learning both semantic and background knowledge of a domain, both of which are required for effective classification on the unlabeled target domain. We also discover that high masking rates are harmful to only the target domain, shedding new light on prior studies that study masking rates in single-domain setups. We hope our study will inspire future investigations in the prompt-based UDA setting.
Ethical Considerations
Our project aims to improve the reliability and safety of language models, which can be fragile under distribution shift Ribeiro et al. (2020) and incur great costs over incorrect predictions Ulmer et al. (2020); Zhang et al. (2021a). By improving performance over distributions without access to labelled data, our method can lead to direct benefits in a wide array of real world applications.
Our study does not involve any human subjects or violation of legal compliance. We do not anticipate any potentially harmful consequences to our work. As detailed in Appendix C, all of our experiments are conducted using publicly available datasets. Our code shall be released for reproducibility. Through our study and releasing our code, we hope to raise stronger research and societal awareness toward the problem of unsupervised domain adaptation in natural language processing.
Limitations and Risks
@@
In our study, we consider a class of PEFT methods that involve inserting learnable parameters between the layers of the model. Other classes of PEFT methods were not considered. However, we use Adapters and He et al. (2022) have shown connections between the method with Prefix Tuning  Li and Liang (2021) and LoRA Hu et al. (2022).
@@
Due to the high variance across runs in PEFT-based learning, we note that the performance can vary significantly across random seeds. We attempt to make our findings reproducible by averaging every experiment over 3 seeds. Taking environmental costs into consideration, we reduce our computational budget by running a majority of our experiments with a smaller-sized model. Learning with larger models is discussed in Appendix F.
Appendix A PEFT Frameworks
The framework proposed in Section 4 is generic, and can be applied to full-model fine-tuning. However, we additionally explore learning in a parameter-efficient setup. Specifically, we use two instantiations in our implementations: Adapters Houlsby et al. (2019) and (IA)${}^{3}$ Liu et al. (2022).
(IA)${}^{3}$
is a state of the art PEFT learning method, and uses around a tenth of learnable parameters compared to popular methods like Adapters. (IA)${}^{3}$ works by element-wise multiplication (i.e. rescaling) of the model’s activations against a learned vector. In this case, the set of learnable parameters $\phi$ is a set of vectors $\{l_{\text{v}},l_{\text{k}},l_{\text{ff}}\}$ applied to each attention mechanism and feed-forward layer as,
Adapters
are a popularly used and high performing PEFT framework, and He et al. (2022) have shown equivalence in the operations applied by Adapters, Prefix Tuning Li and Liang (2021) and LoRA Hu et al. (2022).
$h=h+\gamma(hW_{\text{down}})W_{\text{up}}$
Appendix B Single Phase MLM Training
Our proposed approach in Section 4 involves two stages of training, which is more expensive than standard single phase UDA approaches. In this section, we propose a single training phase variant to FEUDA, and show that it performs similarly to the original method. We use the two phase pipeline in our experiments in the main paper, but note that the single and two phase pipelines are interchangeable.
$(\lambda\;l(\mathbb{C}(x,y);\theta)$
$+(1-\lambda)\;l(\mathbb{M}(x^{\prime});\theta))$
Table 6 compares the performance of the single phase and two phase variants of FEUDA. We also compare with a vanilla joint single phase objective, where $\lambda$ is fixed at 0.5 through training (called Single Phase Vanilla). The performance of the single and two phase variants are almost identical, and either can be used interchangeably. In comparison, the vanilla single phase method is significantly weaker on the target domain.
Method Accuracy
Source Target
Two Phase 93.7 (0.3) 83.3 (0.9)
Singe Phase 93.5 (0.4) 83.3 (0.5)
Singe Phase Vanilla 93.6 (0.1) 75.0 (5.7)
Table 6: Comparison of single and two-phase variants of FEUDA, on the Apparel $\to$ Movies domain pair. The single and two phase variants are almost identical in performance.
Appendix C Preparation of Evaluation Benchmarks
@@
We use two classification datasets, with 5 domains each. This results in a total of 40 pairs of source and target domains. For brevity, we include results of 24 domain pairs in the main paper, and the remaining 16 in Appendix E. For both datasets, we use the train, validation and test splits from Malik et al. (2023). More statistics about each dataset is available in Table 7. The listed datasets are intended for research purposes only. We do not make any commercial use of them.
@@
Appendix D Details on Implementation
@@
We use T5v1.1, T0 and GPT-2 from the HuggingFace library [^4] https://github.com/huggingface/transformers, and use PyTorch Lightning [^5] https://lightning.ai/docs/pytorch/latest/ to train our models. We use the codebase of Liu et al. (2022) [^6] https://github.com/r-three/t-few for implementations of PEFT methods.
@@
Using the (IA)${}^{3}$ PEFT framework, training the T5v1.1 Base model (60 million parameters) for 15,000 steps takes approximately two hours on a single NVIDIA RTX A6000 GPU. The T5v1.1 XL model and T0 model (3 billion parameters) take approximately 8 hours for 15,000 steps of training. For reproducibility, each experiment is repeated thrice, with changing random seeds. In total, we run 540 experiments with the Base model and 72 experiments with the larger models. This results in a total compute time of approximately 2400 GPU hours.
Appendix E Results with Amazon Dataset
Source normal-$\rightarrow$ Target Source Accuracy Target Accuracy
Src Only Src+Tgt MMD FEUDA Src Only Src+Tgt MMD FEUDA
A->B 93.7 (0.1) 93.8 (0.3) 94.3 (0.2) 93.1 (0.3) 93.3 (0.4) 94.7 (0.2) 93.8 (0.3) 93.9 (0.3)
A->B 93.7 (0.1) 94.2 (0.1) 93.8 (0.1) 92.8 (0.6) 90.8 (0.6) 94.3 (0.4) 92.5 (1.1) 90.2 (1.2)
A->C 93.7 (0.1) 93.4 (0.3) 95.0 (0.0) 93.9 (0.5) 91.9 (0.1) 95.0 (0.2) 91.8 (0.5) 92.1 (0.5)
A->M 93.7 (0.1) 94.1 (0.3) 94.7 (0.3) 93.5 (0.4) 81.3 (1.4) 85.8 (0.5) 81.3 (0.6) 83.3 (0.5)
B->A 95.5 (0.2) 94.8 (0.1) 95.8 (0.5) 95.2 (0.2) 93.0 (0.4) 93.4 (0.3) 93.3 (0.2) 93.4 (0.4)
B->Bo 95.5 (0.2) 94.9 (0.1) 95.8 (0.2) 94.3 (0.3) 93.0 (0.9) 94.7 (0.7) 93.8 (0.3) 92.2 (0.1)
B->C 95.5 (0.2) 95.2 (0.2) 96.0 (0.8) 94.6 (0.7) 93.1 (0.3) 94.7 (0.8) 93.4 (0.1) 92.1 (0.3)
B->M 95.5 (0.2) 94.5 (0.4) 96.0 (0.3) 93.9 (0.7) 82.0 (0.5) 85.3 (0.2) 81.3 (0.7) 82.8 (0.2)
Bo->A 94.3 (0.6) 94.7 (0.3) 93.8 (0.4) 91.9 (0.5) 92.3 (0.4) 94.6 (0.3) 91.6 (0.5) 91.3 (0.2)
Bo->B 94.3 (0.6) 94.4 (0.5) 94.9 (0.7) 92.2 (0.3) 93.6 (0.3) 94.8 (0.2) 92.9 (0.6) 90.9 (0.2)
Bo->C 94.3 (0.6) 93.8 (0.7) 94.7 (0.5) 91.6 (0.3) 89.3 (1.1) 94.3 (0.2) 89.8 (0.1) 90.3 (0.4)
Bo->M 94.3 (0.6) 94.3 (0.3) 94.2 (0.4) 91.3 (0.4) 81.7 (0.8) 85.5 (0.9) 84.6 (0.7) 80.1 (1.2)
C->A 93.9 (0.2) 94.6 (0.1) 93.5 (0.2) 93.4 (0.5) 92.1 (0.3) 93.4 (0.4) 92.3 (0.3) 92.5 (0.6)
C->B 93.9 (0.2) 94.3 (0.9) 93.3 (0.7) 93.3 (0.6) 94.0 (0.2) 95.0 (0.6) 94.1 (0.1) 92.1 (0.2)
C->Bo 93.9 (0.2) 93.9 (0.1) 62.3 (0.0) 92.3 (0.2) 91.1 (0.4) 93.9 (0.8) 91.3 (0.5) 89.0 (0.1)
C->M 93.9 (0.2) 94.1 (0.3) 92.9 (0.9) 93.4 (0.5) 82.3 (1.1) 85.8 (0.1) 81.5 (0.7) 79.7 (1.2)
M->A 85.5 (0.2) 85.6 (0.7) 86.3 (0.6) 83.3 (0.6) 89.8 (0.8) 94.2 (0.7) 89.1 (1.4) 90.1 (0.5)
M->B 85.5 (0.2) 86.1 (0.4) 78.3 (11.1) 83.7 (0.3) 91.7 (0.5) 95.3 (0.5) 81.0 (16.1) 89.9 (1.2)
M->Bo 85.5 (0.2) 84.6 (0.8) 76.4 (13.7) 83.8 (0.0) 92.7 (0.1) 94.1 (0.4) 80.5 (18.6) 91.5 (0.0)
M->C 85.5 (0.2) 86.1 (0.7) 87.0 (0.0) 84.6 (0.6) 90.1 (0.3) 94.3 (0.5) 90.5 (0.0) 89.7 (0.3)
Table 8: Comparison of FEUDA and MMD by classification accuracy on the Amazon Product Review dataset, using the T5v1.1 base model, and (IA)${}^{3}$ PEFT method. FEUDA is competitive with MMD on average. For the target domain, the highest values between FEUDA and MMD have been marked in bold. Cases where Src Only outperforms both FEUDA and MMD have been underlined.
Figure 4: UMap visualizations of sentence embeddings from the Apparel $\to$ Movies data pair, using the T5v1.1 base model and (IA)${}^{3}$ PEFT method. Despite not promoting domain-invariance, FEUDA learns sentence embeddings that are separable by class labels, regardless of the domain of these sentences. The classification hyperplane for the source domain has been imagined as a solid line for illustration purposes, and its extension to the target domain is shown as a dashed line.
Table 8 shows the performance of our proposed approach on the Amazon Product Reviews dataset. On average, FEUDA is competitive with the state of the art MMD method from Malik et al. (2023).
@@ Appendix E Results with Amazon Dataset
We confirm this by checking for a significant difference in the performance of FEUDA and MMD on the 20 dataset pairs. The Mann-Whitney U test and Student’s t-test both resulted in non-significant p-values of 0.5516 and 0.8316, confirming the hypothesis that there is no significant difference between FEUDA and MMD on the Amazon dataset. However, on the MNLI dataset, where all domains have larger gaps, both significant tests showed a significant difference between FEUDA and MMD, with FEUDA being more powerful.
FEUDA can learn representations that generalize across domains
@@
Appendix F Learning with Larger Models
The use of high capacity language models, which are too compute intensive for traditional fine-tuning, originally encouraged the use of prompting, and prompting is now common with learning from large language models Brown et al. (2020); Schick and Schütze (2021a, b); Gao et al. (2021). Since our approach makes use of prompting, we investigate the performance of FEUDA with such models.
We experiment with two large models: T5v1.1 XL (3 billion parameters) and T0 (3 billion parameters). T0 is optimized for zero-shot generalization by training on supervised prompts. We use the Apparel (A) $\to$ Movies (M) domain pair from the Amazon Reviews dataset. Table 9 shows the performance gap between FEUDA and MMD increasing with larger models. In the case of T0, we see particularly poor performance with MMD. This may be due to the fact that the task of minimizing the divergence between embeddings is highly different from the tasks a model is trained on during instruction tuning.
Model Src Only Src+Tgt MMD FEUDA
Source Target Source Target Source Target Source Target
T5 v1.1 Base 93.7 (0.1) 81.3 (1.4) 94.1 (0.3) 85.8 (0.5) 93.4 (0.2) 78.6 (1.3) 93.5 (0.4) 83.3 (0.5)
T5 v1.1 XL 95.4 (0.2) 89.1 (0.8) 95.3 (0.6) 93.0 (0.5) 74.8 (15.6) 65.2 (9.5) 95.1 (0.2) 92.0 (1.5)
T0 3B 95.5 (0.4) 91.3 (0.2) 95.5 (0.2) 92.2 (0.7) 52.1 (1.1) 51.8 (0.8) 95.5 (0.4) 93.8 (0.4)
Table 9: Performance of FEUDA across the T5v1.1 Base (60 million parameters), T5v1.1 XL (3 billion parameters) and T0 (3 billion parameters) models on the Apparel $\to$ Movies domain pair. We report the mean and standard deviation over 3 runs. The performance gap between FEUDA and MMD increases with larger models.
Appendix G More On Analyzing the Impact of MLM on UDA
Table 10 accompanies the results from Figure 1, which shows the impact of masking sequences at inference. Words are selected for masking based on their their “informativeness”, measured by their PMI to the inference class label. The performance of the model is best with the original unmasked sequences, indicating the presence of both informative and uninformative words are essential for strong classification performance.
To isolate any effects of PEFT methods or pre-training data, we repeat the analysis from Table 3 with fine-tuning Flan-T5 and notice a similar trend in Table 11.
Table 12 accompanies results from Figure 2, which show the impact of varying masking rates on FEUDA. Using the T5v1.1 base model, we train FEUDA using varying random masking rates on the Apparel $\to$ Movies domain pair, and report the mean and standard deviation over three runs. With high masking rates, the performance on the source domain is largely maintained, but the performance on the target domain rapidly deteriorates.
@@ Appendix G More On Analyzing the Impact of MLM on UDA
Source Target
Original 93.5 83.3
Informative Masking 88.0 78.8
Uninformative Masking 92.0 79.0
Table 10: Impact of masking at inference. We evaluate FEUDA on the Apparel $\to$ Movies domain pair, and select words for masking based on their “informativeness” to the classification task. The performance of the model is best with the original unmasked sequences, indicating the presence of both informative and uninformative words are essential for strong classification performance.
Masking Strategy Accuracy
Source Target
Random 95.8 (0.0) 86.8 (0.3)
Informative 93.9 (0.6) 85.3 (0.3)
Uninformative 95.0 (0.0) 84.8 (0.1)
Table 11: Impact of word selection for masking during training. Using Flan-T5 base and no PEFT methods, we find that masking words at random is more powerful than selectively masking informative or uninformative words. This indicates that the model requires both semantic features (learnt through masking informative words) and background features (learnt through masking uninformative words) for classification on the unlabelled target domain.

### v2

@@ 9 Conclusion
We introduce the setting of Generative UDA, and perform an investigation on the utility of continued pre-training in this setting. We compare the approach with the popular class of domain-invariance based methods for UDA, showing that CPT is both competitive with, and more stable than invariance based approaches. Our experiments show that the benefits of CPT extend to different architectures, tuning methods and data regimes. We motivate the need for target domain exposure through CPT by showing that performance on the target domain gradually degrades with increasing masking rate. Finally, we shed light on the interplay between masking and classification performance, and how this aids UDA. Our analysis shows that in predicting masked words that are informative to the downstream task, the model implicitly learns about the downstream task, furthering the benefits of directly learning the task. Our work connects the body of UDA research with that of instruction tuning, enabling an initial step towards a wider applicability of modern language models.
Limitations
Our work presents an investigation into continued pre-training for UDA in a generative setting. Since generative UDA is an almost completely unexplored area, we establish a proof of concept by using sentence classification tasks for our analysis. We leave the extending our analysis to more complex tasks to future work.
@@
In our study, we consider a class of PEFT methods that involve inserting learnable parameters between the layers of the model. Other classes of PEFT methods were not considered. However, we use Adapters and He et al. (2022) have shown connections between the method with Prefix Tuning  (Li & Liang, 2021) and LoRA (Hu et al., 2022).
@@
Due to the high variance across runs in PEFT-based learning, we note that the performance can vary significantly across random seeds. We attempt to make our findings reproducible by averaging every experiment over 3 seeds. Taking environmental costs into consideration, we reduce our computational budget by running a majority of our experiments with a smaller-sized model. Learning with larger models is discussed in Section 4.
Ethics Statement
Our project aims to extend the problem of unsupervised domain adaptation to the generative setting, matching current needs with large language models. This is an effort towards improving the reliability and safety of language models, which can be fragile under distribution shift Ribeiro et al. (2020) and incur great costs over incorrect predictions Ulmer et al. (2020); Zhang et al. (2021a).
Our study does not involve any human subjects or violation of legal compliance. We do not anticipate any potentially harmful consequences to our work. As detailed in Appendix A, all of our experiments are conducted using publicly available datasets. Our code shall be released for reproducibility. Through our study and releasing our code, we hope to raise stronger research and societal awareness toward the problem of unsupervised domain adaptation in natural language processing.
Appendix A Preparation of Evaluation Benchmarks
@@
We use two classification datasets, with 5 domains each. This results in a total of 40 pairs of source and target domains. For brevity, we include results of 24 domain pairs in the main paper, and the remaining 16 in Appendix C. For both datasets, we use the train, validation and test splits from Malik et al. (2023). More statistics about each dataset is available in Table 4. The listed datasets are intended for research purposes only. We do not make any commercial use of them.
@@
Appendix B Details on Implementation
@@
We use T5v1.1, T0 and GPT-2 and LLaMA-2 from the HuggingFace library [^8] https://github.com/huggingface/transformers, and use PyTorch [^9] https://pytorch.org/ to train our models.
@@
[…]
Training Steps Source Accuracy Target Accuracy
5,000 93.2 (0.4) 81.9 (0.4)
10,000 93.4 (0.5) 81.6 (0.6)
15,000 93.5 (0.4) 83.3 (0.5)
[…]
@@
Using the (IA)3 PEFT framework, training the T5v1.1 Base model (60 million parameters) for 15,000 steps takes approximately two hours on a single NVIDIA RTX A6000 GPU. The T5v1.1 XL model and T0 model (3 billion parameters) take approximately 8 hours for 15,000 steps of training. For reproducibility, each experiment is repeated thrice, with changing random seeds. In total, we run 540 experiments with the Base model and 72 experiments with the larger models. This results in a total compute time of approximately 2400 GPU hours.
Appendix C Detailed results with the Amazon and MNLI Datasets
Table 6 shows the performance of CPT on the Amazon and MNLI datasets.
@@ Appendix C Detailed results with the Amazon and MNLI Datasets
On the Amazon dataset, CPT is competitive with the state of the art UDAPTER method from Malik et al. (2023) on average. We confirm this by checking for a significant difference in the performance of CPT and UDAPTER on the 20 dataset pairs. The Mann-Whitney U test and Student’s t-test both resulted in non-significant p-values of 0.5516 and 0.8316, confirming the hypothesis that there is no significant difference between CPT and UDAPTER on the Amazon dataset.
However, on the MNLI dataset, where all domains have larger gaps, both significant tests showed a significant difference between CPT and UDAPTER, with CPT being more powerful. This is exemplified through cases like Travel (T) $\to$ Government (G), where CPT yields an accuracy of 83.6% on the target domain, equalling the upper bound of the Src+Tgt baseline.
Amazon MNLI
Src+Tgt UDAPTER CPT Src+Tgt UDAPTER CPT
A $\rightarrow$ B 94.7 (0.2) 93.8 (0.3) 93.9 (0.3) T $\rightarrow$ F 77.2 (0.4) 69.7 (0.8) 74.1 (0.9)
A $\rightarrow$ Bo 94.3 (0.4) 92.5 (1.1) 90.2 (1.2) T $\rightarrow$ G 83.6 (0.7) 79.3 (0.5) 83.6 (0.3)
A $\rightarrow$ C 95.0 (0.2) 91.8 (0.5) 92.1 (0.5) T $\rightarrow$ S 72.3 (0.5) 69.6 (0.1) 70.7 (0.6)
A $\rightarrow$ M 85.8 (0.5) 81.3 (0.6) 83.3 (0.5) T $\rightarrow$ Te 77.8 (0.1) 69.4 (0.8) 76.8 (0.0)
B $\rightarrow$ A 93.4 (0.3) 93.3 (0.2) 93.4 (0.4) F $\rightarrow$ T 79.9 (0.1) 69.9 (0.2) 65.4 (0.8)
B $\rightarrow$ Bo 94.7 (0.7) 93.8 (0.3) 92.2 (0.1) F $\rightarrow$ G 82.3 (0.1) 54.3 (23.4) 78.8 (2.5)
B $\rightarrow$ C 94.7 (0.8) 93.4 (0.1) 92.1 (0.3) F $\rightarrow$ S 72.1 (0.2) 64.6 (1.8) 65.3 (1.6)
B $\rightarrow$ M 85.3 (0.2) 81.3 (0.7) 82.8 (0.2) F $\rightarrow$ Te 78.3 (0.6) 64.6 (0.7) 72.5 (0.2)
Bo $\rightarrow$ A 94.6 (0.3) 91.6 (0.5) 91.3 (0.2) G $\rightarrow$ T 79.9 (0.4) 75.9 (0.3) 75.8 (0.6)
Bo $\rightarrow$ B 94.8 (0.2) 92.9 (0.6) 90.9 (0.2) G $\rightarrow$ F 76.7 (0.1) 69.9 (0.2) 73.5 (0.2)
Bo $\rightarrow$ C 94.3 (0.2) 89.8 (0.1) 90.3 (0.4) G $\rightarrow$ S 73.1 (0.0) 69.4 (0.1) 68.0 (1.8)
Bo $\rightarrow$ M 85.5 (0.9) 84.6 (0.7) 80.1 (1.2) G $\rightarrow$ Te 78.1 (0.6) 69.9 (0.3) 73.5 (0.6)
C $\rightarrow$ A 93.4 (0.4) 92.3 (0.3) 92.5 (0.6) S $\rightarrow$ T 79.5 (0.3) 74.4 (1.7) 76.8 (0.1)
C $\rightarrow$ B 95.0 (0.6) 94.1 (0.1) 92.1 (0.2) S $\rightarrow$ F 77.7 (0.2) 73.1 (0.0) 72.4 (0.5)
C $\rightarrow$ Bo 93.9 (0.8) 91.3 (0.5) 89.0 (0.1) S $\rightarrow$ G 83.4 (0.2) 78.2 (0.5) 76.3 (0.9)
C $\rightarrow$ M 85.8 (0.1) 81.5 (0.7) 79.7 (1.2) S $\rightarrow$ Te 78.5 (0.0) 66.7 (0.2) 74.8 (0.3)
M $\rightarrow$ A 94.2 (0.7) 89.1 (1.4) 90.1 (0.5) Te $\rightarrow$ T 79.8 (0.3) 71.4 (0.0) 76.5 (0.4)
M $\rightarrow$ B 95.3 (0.5) 81.0 (16.1) 89.9 (1.2) Te $\rightarrow$ F 77.9 (0.1) 69.9 (0.5) 74.3 (0.5)
M $\rightarrow$ Bo 94.1 (0.4) 80.5 (18.6) 91.5 (0.0) Te $\rightarrow$ G 82.5 (0.1) 75.6 (1.6) 82.0 (0.6)
M $\rightarrow$ C 94.3 (0.5) 90.5 (0.0) 89.7 (0.3) Te $\rightarrow$ S 72.2 (0.0) 68.0 (0.4) 71.3 (0.5)
@@
Figure 8: UMap visualizations of sentence embeddings from the Apparel $\to$ Movies data pair, using the T5v1.1 base model and (IA)3 PEFT method. Despite not promoting domain-invariance, CPT may be learning sentence embeddings that are separable by class labels, regardless of the domain of these sentences. The classification hyperplane for the source domain has been imagined as a solid line for illustration purposes, and its extension to the target domain is shown as a dashed line.
@@
CPT 83.3 (0.9)
UDAPTER 81.3 (0.6)
DANN 52.3 (1.7)
CORAL 80.9 (0.4)
Task Vectors 48.0 (0.7)
Task Vectors (fine-tuning) 69.0 (0.4)
[…]
Comparison with other Model Centric Approaches
In addition to the MMD based method of Malik et al. (2023), we also compare CPT with other methods that promote domain invariance: 1) DANN (Ganin et al., 2016), which is the most widely used UDA method in NLP (Ramponi & Plank, 2020), but has been shown to be highly unstable; 2) CORAL (Sun et al., 2017), which minimizes second order statistics of the data embeddings. Additionally, with an emerging class of weight interpolation based methods, we make a comparison with task vector arithmetic Ilharco et al. (2022). The use of task vectors with PEFT methods beyond LoRA (Hu et al., 2022) has been unexplored in the literature, and we find that the method does not work with IA3. With fully fine-tuned models, the method improves in performance, but is still weaker than CPT.
CPT may learn representations that generalize across domains
[…]
Appendix D CPT across Model Architectures and Scales
We evaluate the performance of CPT over T5v1.1 XL and the instruction tuned T0 (3B) (Sanh et al., 2022) in Table 8.
Model Src+Tgt UDAPTER CPT
T5 v1.1 Base 85.8 (0.5) 78.6 (1.3) 83.3 (0.5)
T5 v1.1 XL 93.0 (0.5) 65.2 (9.5) 92.0 (1.5)
T0 3B 92.2 (0.7) 51.8 (0.8) 93.8 (0.4)
Table 8: The performance gap between CPT and UDAPTER increases with larger models, from T5v1.1 Base (60M parameters) to T5v1.1 XL (3B parameters), and further increases with instruction tuning (T0 3B).
Appendix E PEFT Frameworks
The framework proposed in Section 2 is general and can be applied to fine-tune all model parameters. Additionally, our CPT framework is compatible with the parameter-efficient fine-tuning approach. The PEFT approach is desirable because it adds only a small amount of learnable parameters $\phi$ to a pre-trained language model $\theta$, and fine-tunes only $\phi$ to perform prediction while keeping the other model parameters $\theta$ frozen. We use two instantiations in our implementations: Adapters Houlsby et al. (2019) and (IA)3 Liu et al. (2022).
(IA)3
is a state of the art PEFT learning method, and uses around a tenth of learnable parameters compared to popular methods like Adapters. (IA)3 works by element-wise multiplication (i.e. rescaling) of the model’s activations against a learned vector. In this case, the set of learnable parameters $\phi$ is a set of vectors $\{l_{\text{v}},l_{\text{k}},l_{\text{ff}}\}$ applied to each attention mechanism and feed-forward layer as,
Adapters
are a popularly used and high performing PEFT framework, and He et al. (2022) have shown equivalence in the operations applied by Adapters, Prefix Tuning Li & Liang (2021) and LoRA Hu et al. (2022).
$h=h+\gamma(hW_{\text{down}})W_{\text{up}}$
Table 9 shows CPT beats UDAPTER across different tuning methods. We also note that fine-tuning yields slightly better performance for all UDA methods.
Method Src+Tgt UDAPTER CPT
Fine-Tuning 86.4 (0.4) 82.4 (1.6) 84.4 (0.3)
(IA)3 85.8 (0.5) 81.3 (0.6) 83.3 (0.5)
Adapters 85.3 (0.5) 79.1 (0.3) 82.7 (0.5)
Table 9: Performance of CPT across different adaptation methods with the T5v1.1 base model on the Apparel $\to$ Movies domain pair. CPT remains more powerful than UDAPTER across all methods.
Appendix F CPT in a Few-Shot Setup
Table 10 accompanies Figure 2 (Section 4), showing the 256-shot performance of CPT and other baselines, across model sizes. Similarly, Table 11 accompanies Figure 4, showing the relative performance of all baselines across varying $k$.
Model Src+Tgt UDAPTER CPT
T5v1.1 Base 77.8 (0.4) 60.9 (1.6) 73.1 (1.7)
T5v1.1 XL 84.4 (0.1) 84.8 (1.5) 89.9 (1.1)
T0 3B 88.3 (0.5) 81.8 (1.3) 93.9 (0.4)
Model Src+Tgt UDAPTER Two Phase CPT
T5v1.1 Base 82.8 (0.6) 62.5 (0.7) 79.8 (1.4) 81.2 (0.7)
T5v1.1 XL 92.5 (0.4) 71.7 (7.8) 84.3 (0.9) 86.8 (2.2)
T0 3B 91.8 (0.6) 79.5 (6.7) 53.5 (0.4) 92.8 (0.2)
Table 10: Performance of CPT across different models, in a k-shot learning setup on the Apparel $\to$ Movies domain pair. We see CPT retaining strong performance on the target domain across models. Left: 32-shot. Right: 256-shot.
Number of Shots Src+Tgt UDAPTER Two Phase UDAPTER CPT
32 77.8 (0.4) 60.9 (1.6) 59.4 (2.0) 73.1 (1.7)
128 82.5 (0.5) 75.1 (0.6) 62.8 (0.6) 78.8 (1.0)
256 82.8 (0.6) 62.5 (0.7) 79.8 (1.4) 81.2 (0.7)
Table 11: Performance of CPT across different number of shots, on the Apparel $\to$ Movies domain pair, using the T5v1.1 base model. We see CPT retaining strong performance on the target domain across shots.
Appendix G Impact of Target Domain Exposure
[…]
Table 12 accompanies results from Figure 5, which show the impact of varying masking rates on CPT. Using the T5v1.1 base model, we train CPT using varying random masking rates on the Apparel $\to$ Movies domain pair, and report the mean and standard deviation over three runs. With high masking rates, the performance on the source domain is largely maintained, but the performance on the target domain rapidly deteriorates.